For a home repair job you must turn the handle of a screwdriver 32 times.

Part A

If you apply an average force of 15 N tangentially to the 2.0-cm-diameter handle, how much work have you done?

Part B

If you complete the job in 22 seconds, what was your average power output?
Answer #1

n = 32 times

(A)
F = 15 N
r = 2.0/200 = 0.01 m

Work done = F.d
Where d is the distance moved , = 32 * 2*π*r

W = 15 * 32* 2 * π * 0.01
W =  30.16 J


(b)
t = 22 s
Power = Word done / time
P = 30.16/22
P = 1.37 W